Physics 031-Exam page 3 PART IL You must justify y
Answer #1

onl 22-34 unit/,- 4.47) 냐.YFㄴ 2. 3

m2g - T1 =m2.a ....... eqn1
T1 - T2 = m1.a ...... eqn2

from 1
T1 =m2 (g-a)

from 2
T2= m2(g-a)-m1.a =(m2 - m1)a-m2g

so T2/T1 = (m2 - m1)a-m2g/m2 (g-a)= ma-2mg/2m(g-a) = (a-2g)/2(g-a)